### v1

@@ Abstract
Coherent elastic neutrino-nucleus scattering (CENS) provides a unique probe for neutrino properties Beyond the Standard Model (BSM) physics. REactor neutrino LIquid xenon Coherent Scattering experiment (RELICS), a proposed reactor neutrino program using liquid xenon time projection chamber (LXeTPC) technology, aims to investigate the CENS process of antineutrinos off xenon atomic nuclei. In this work, the design of the experiment is studied and optimized based on Monte Carlo (MC) simulations. To achieve a sufficiently low energy threshold for CENS detection, an ionization-only analysis channel will be adopted for RELICS. A high emission rate of delayed electrons after a big ionization signal is the major background, leading to an analysis threshold of 120 photo-electrons in the CENS search. The second largest background, nuclear recoils induced by cosmic-ray neutrons, is suppressed via a passive water shield. The physics potential of RELICS was explored with a 32 exposure at a baseline of 25 m from a reactor core with a 3 GW thermal power. In an energy range of 120 to 240 PEs, we the expected 13673.5 CENS and 2133.7 background events.The sensitivity of RELICS to the weak mixing angle was investigated at a low momentum transfer. Our study has shown that RELICS can further improve the constraints on the non-standard neutrino interaction (NSI) compared to the best results. RELICS set out to develop a powerful surface-level detection technology for low-energy neutrinos from reactors.
@@
Coherent elastic neutrino-nucleus scattering (CENS) is a Standard Model (SM) weak neutral current process first proposed in 1974 [1]. The neutrino treats the nucleus as a single particle during scattering with a cross-section proportional to the nucleon number squared. Despite this process having a relatively large cross-section, the detection is challenging due to the ultra-low deposited energy. This process had evaded experimental discoveries for decades until 2017 when it was first observed by COHERENT collaboration with neutrinos produced by the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory scattering off a CsI scintillator [2], and later confirmed with a liquid argon detector [3] in 2020. These measurements of CENS offer new probes for diverse physical phenomena.
CENS can provide new constraints on physics beyond the Standard Model (SM). The weak mixing angle is related to the CENS cross-section [4]. Any deviation from the SM prediction in measurement can be an indication of new physics. CENS is also a valuable tool to constrain nonstandard neutrino interactions (NSI) [5], to search for sterile neutrinos [6], to verify the existence of non-trivial neutrino electromagnetic properties [7], and to study the nuclear structure [8] and astrophysics [9]. In addition, the CENS signal from the solar or atmospheric neutrinos is expected to be an irreducible background for the next-generation dark matter direct detection experiments, such as PandaX-30T [10] and XLZD [11]. A precise measurement of this process can provide valuable information for dark matter experiments digging into the so-called “neutrino floor/fog” [12, 13].
@@
RELICS is a reactor neutrino experiment using the LXeTPC technology. LXeTPC is one of the leading technologies for rare event searches, such as detecting dark matter and neutrinoless double beta decay [11, 25, 26, 27, 28, 29]. Benefiting from its ultra-low threshold by using ionization-only channel [30, 31] and reduced background, LXeTPC is an excellent approach for reactor CENS detection. The RELICS experiment is proposed to be located near the Sanmen Nuclear Power Plant in Zhejiang province, China. At a 25 m baseline to the 3 GW reactor core, the average neutrino flux available for RELICS is 1013cm-2s-1.
@@ IV Expected CENS events
where is Fermi’s coupling constant, is the weak nuclear charge, is the mass of the target nucleus, is the ground state elastic form factor, is the nuclear recoil energy while the incident neutrino energy. is required to ensure coherence, where is the nuclear radius. Normally, this can be satisfied when MeV.
@@ V Background contributions
single-scatter: Given the small interaction cross-section of neutrino, the probability of a neutrino having more than one interaction in the sensitive LXe volume is negligible. We require only one energy deposition above 0.5 keVnr energy in the LXe volume.
@@
In addition, dedicated selection methods are developed to suppress backgrounds out of the delayed electrons, such as waveform classifier and pattern classifier. These two criteria are described in detail in (V.5).
@@ V.1 Background induced by cosmic-ray neutrons
The cosmic rays that travel through the atmosphere can produce a variety of radiation particles, such as neutrons, protons, pions, etc, through spallation reaction on nitrogen, oxygen, etc [48, 49]. Among them, the high-energy cosmic-ray neutrons (CRNs) can easily pass through the shield and deposit energy in the RELICS detector. The main purpose of the water shield design is to block the CRNs. The size of the water tank is optimized based on the MC simulation. The CRN flux is calculated using the Cosmic-Ray Shower Generator (CRY) source [50]. The energy spectrum of the CRNs from CRY is shown in figure 6. Analyze the MC data shows that only CRNs with energy greater than 103 MeV can deposit energy in LXe, which also indicates that the neutrons from the nuclear reactor and laboratory environment normally with energy less than 10 MeV [51] that also shows in figure 6, can not contribute to the backgrounds.
After processing the MC data, and passing the LXe-veto, single-scatter and FV selections, the CRNs-induced NR background rate is estimated to be (7.7 0.7) in keVnr region-of-interest (ROI) of CENS detection, with the spectrum shows in figure 7 (top). In addition, the secondaries of CRNs will also induce ER background in LXe, with a rate of (9.0 0.3), which is shown in figure 7 (bottom).
@@ V.2 Background induced by cosmic muons
The RELICS detector will operate without overburden. Hence, special attention has to be paid to the muon-related background. Even though muons are tagged with 99% efficiency by the PS veto detector, the neutron background produced by muon interacting with the detector materials is still one of the main background components. Dedicated simulations are performed to evaluate the muon-induced backgrounds. The energy and angular distributions of muons follow the Shukla model [52]. The energy deposition in LXe that is caused by muon-induced fast neutron and related gammas are recorded by RelicsSim, and then filtered with LXe-veto, single-scatter and FV selections. Assuming a 99% muon veto tagging efficiency, the muon-induced NRs and ERs are (0.08 0.01) in keVnr and (0.66 0.09) in keVer, respectively. The corresponding spectra are shown in figure 7.
@@ V.3 Neutrons and gammas from the detector materials
Materials constructing the detector contain radioactive contamination. The selection of the materials requires an extensive radioactivity screening campaign for rare event search experiments. To evaluate the background contribution from these materials, the radioactive contamination for all the materials is considered in the RELICS MC model. Currently, the radioactivity of isotopes from XENON100 [53] is adopted. A dedicated effort is also being carried out to select clean materials for the RELICS experiment, leading to similar radioactivity for most of the materials.
The NR background rate induced by neutrons is around one event per year and completely negligible in the CENS ROI. ER background induced by gammas is studied in this work. The decay of these isotopes is generated and confined uniformly in the corresponding detector component in the simulation. By applying successive selection criteria, including LXe-veto, single-scatter and FV selections, the spatial distribution of ER background events inside the active LXe volume in the energy region of [0, 40] keVer is shown in figure 8. Figure 7 (bottom) shows the corresponding energy spectrum, which is two orders of magnitude higher than the ER background components. The average ER background rate is (364 5) in the FV volume.
S2 width selection: As shown in figure 8, a large amount of background events are located close to the liquid-gas interface. These backgrounds are mainly from the radioactivities of PMTs and can not be suppressed without the selection in “z” in an “S2-only” analysis. As demonstrated in [30], the pulse shape of S2 signals can discriminate background happening at the liquid-gas interface to signals distributed uniformly in “z”. A toy MC was conducted to consider the drift and diffusion properties of electrons in LXe to optimize such a selection. We simulate S2s and their pulse shape parameters with “z” distributions shown in figure 8 with energy in the CENS ROI. The S2s are simulated via RelicsAPT where the NEST model is incorporated. The S2 pulse shape is simulated using the electron diffusion model , where is the longitudinal diffusion coefficient, is the electron drift velocity, is the width of single electron waveform [54]. In our simulation. = 16 cm2/s, = 0.174 cm/s, and = 0.19s.
@@ V.4.1 Intrinsic 222Rn and 85Kr
Krypton mixes into xenon gas when it is extracted from the air. 85Kr, regardless of its low abundance of in natural krypton, is one of the main backgrounds in any LXe experiments searching for rare events. The concentration of krypton in xenon is usually reduced below ppt level in dark matter search experiments, such as XENONnT and PandaX-4T, through distillations with dedicated columns[ref]. In RELICS, we assume a natKr/Xe concentration of 10 ppt and expect (0.28 0.01)10 as shown in figure 7 (bottom), which is a subdominant component.
The other intrinsic background in LXe comes from the decays of 222Rn, which emanates from any material touching Xenon gas or liquid. The emanated 222Rn diffuses homogeneously inside the LXe volume within its relatively long half-life of 3.8 days. The decay of 214Pb is another important background component in the CENS ROI in the decay chain of 222Rn. With similar size and material, the radioactivity of 222Rn in RELICS is comparable to XENON100. Assuming a conservative 222Rn radioactivity of 40 as achieved by the XENON100 experiment [53] (20 ), the expected ER background rate is estimated as of (0.98 0.02)10-3 kg-1day-1keV-1 as shown in figure 7 (bottom), which is also a subdominant component in the CENS search.
@@ V.4.2 Cosmogenic isotopes and
: Natural xenon does not contain . can be produced from when cosmogenic muons kick off a neutron in the nucleus with a cross-section of (2.74 0.4) cm2. Additionally, we consider another production channel through the neutron capture on with an upper limit cross-section of 0.7cm2 at 95% C.L [56]. The production rate of in LXe is calculated based on its production cross-section, the muon’s fluxes and energy spectrum, and thermal neutrons. The decays to via electron capture (EC) process, with a half-life of 36.4 days. We assume that the production and decay of reach equilibrium before deploying the RELICS detector inside the shielding. The estimated equilibrium decay rate is 9.29 evts/day/kg. The shielding system of RELICS can attenuate neutrons efficiently, and the equilibrium decay rate can decrease to 0.79 evts/day/kg after 300 days.
@@
: We consider three possible mechanisms for production in xenon, following the approaches in [57]. Firstly, the extremely low content of in the atmosphere, which is generated by spallation of through (n, 4n), neutron capture of , or cosmic bombardment of calcium-containing soils, via (n, ). This mechanism only needs to be considered in the context of air leaks of an LXe detector, which is strictly limited to below 0.1 liter per year to ensure the purity level of LXe. Secondly, the spallation of detector materials, especially iron in steel. However, its contribution is greatly suppressed due to the slow diffusion rate of argon in steel. Lastly, the direct spallation of xenon atoms by high-energy cosmogenic protons and neutrons. The semi-empirical formula by Silberberg and Tsao was used to calculate the spallation cross-section. The CRY is used to calculate the proton spectrum. Gordon’s neutron spectrum [58] is adopted for cosmogenic neutron estimation. The production rate of this process is estimated to be 0.024 atoms/kg/day with an uncertainty of 100%. The produced decays to the ground state of via EC and subsequently deposits energy via x-rays or Auger electrons. The total energy (relative intensity) for K, L, and M shells are 2.82 keV (90.2%), 0.270 keV (8.9%), and 0.018 keV (0.9%), respectively. As a results, the background rate of in the CENS ROI is below kg-1day-1 and completely negligible for CENS search.
@@ V.5 Delayed electrons
The first kind of delayed electrons is produced via photon-ionization of impurities or metals in the drift region of the TPC. Photon-ionized delay electrons only appear shortly (within a time range that is comparable to the largest drift time of the LXeTPC) after the high energy S2 signals and will be completely subdominant when a S exposure is rejected for each of the high energy S2s, leading to an exposure loss of 1%.
Besides photon-ionization, delayed electrons can also be produced in a larger time range of up to a few seconds. The rate of delayed electrons is observed to be correlated with the level of impurities in LXe and the amplitude of the extraction field. Muons passing through the LXeTPC result in large energy deposition, generating a substantial number of ionization electrons. These delayed electrons may occur after the primary S2 signal of muon events, leading to a prolonged tail of signals composed of a single or a few electrons. Due to the high rate of cosmic muons, RELICS faces a major challenge in suppressing background from accidental pile-ups of delayed electrons in a short time range that is comparable to the width of a physical S2 signal. This background is hereafter referred to as Delayed Electron (DE).
@@ V.5.1 Expected rate of DE
As shown in previous studies with XENON1T [59] and LUX [60], the emission rate of delayed electrons are correlated, in time and position, with the previous high-energy events. This part of the background is referred to as a correlated delayed electron background and can be significantly suppressed due to its unique correlation with proceeding high-energy events. We simulate muons passing through the TPC active volume using Geant4, tracking their interaction vertex and the energy deposited along the track. The ionization signals are generated along the muon track, assuming a constant ionization yield of 50 electrons per keV. We then simulated the generation of delayed electrons, assuming that 0.1% of the total ionization electrons from primary interactions are delayed emissions in a time range of 0.1 ms to 2 s, comparable to [59]. The distribution in “delayed time” (time respected to the primary S2) for delayed electrons is assumed to follow a power law with a coefficient of -1.1, as shown in XENON1T [59]. These delayed electrons’ horizontal positions () are assumed to be the same as their primary S2s. Position reconstruction resolution for single electron S2s, which is expected to be 20 mm from optical simulations, is taken into account in this analysis. Ionization S2 signals are grouped together if the time gap between them is below 1 s in order not to split S2 waveforms of events happening at the bottom part of the TPC. As a result, the pile-up of DE (called Pile-up DEs) leads to a significant background in CENS search. In RELICS, we expect more than 10 Hz of muon events in the LXeTPC, leaving a large amount of single electron S2 signals. The expected event rate of a single-electron pile-up background is summarized in Table 2. Without implementing appropriate cuts to reduce the rate of these pile-up events, CENS signals with 3 to 6 electrons will be completely overwhelmed.

### v2

@@ Abstract
Coherent elastic neutrino-nucleus scattering (CENS) provides a unique probe for neutrino properties Beyond the Standard Model (BSM) physics. REactor neutrino LIquid xenon Coherent Scattering experiment (RELICS), a proposed reactor neutrino program using liquid xenon time projection chamber (LXeTPC) technology, aims to investigate the CENS process of antineutrinos off xenon atomic nuclei. In this work, the design of the experiment is studied and optimized based on Monte Carlo (MC) simulations. To achieve a sufficiently low energy threshold for CENS detection, an ionization-only analysis channel is adopted for RELICS. A high emission rate of delayed electrons after a big ionization signal is the major background, leading to an analysis threshold of 120 photo-electrons in the CENS search. The second largest background, nuclear recoils induced by cosmic-ray neutrons, is suppressed via a passive water shield. The physics potential of RELICS is explored with a 32 exposure at a baseline of 25 m from a reactor core with a 3 GW thermal power. In an energy range of 120 to 240 PE, we expect 4902.4 CENS and 1318.4 background events. The sensitivity of RELICS to the weak mixing angle is investigated at a low momentum transfer. Our study shows that RELICS can further improve the constraints on the non-standard neutrino interaction (NSI) compared to the current best results.
@@
Coherent elastic neutrino-nucleus scattering (CENS) is a Standard Model (SM) weak neutral current process proposed in 1974 [1]. The neutrino treats the nucleus as a single particle during scattering with a cross-section proportional to the nucleon number squared. Despite this process having a relatively large cross-section, the detection is challenging due to the ultra-low deposited energy. This process had evaded experimental discoveries for decades until 2017 when it was first observed by the COHERENT collaboration with neutrinos produced by the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory scattering off a CsI scintillator [2], and later confirmed with a liquid argon detector [3] in 2020. These measurements of CENS offer new probes for diverse physical phenomena.
CENS can provide new constraints on the Standard Model physics and the physics Beyond it (BSM). The weak mixing angle is related to the CENS cross-section [4]. Any deviation from the SM prediction in measurement can be an indication of new physics. CENS is also a valuable tool to constrain nonstandard neutrino interactions (NSI) [5], to search for sterile neutrinos [6], to verify the existence of non-trivial neutrino electromagnetic properties [7], to study the nuclear structure [8] and astrophysics [9]. In addition, the CENS signal from the solar or atmospheric neutrinos is expected to be an irreducible background for the next-generation dark matter direct detection experiments, such as PandaX-xT [10] and XLZD [11]. A precise measurement of this process, especially one with a xenon target, can provide valuable information for dark matter experiments digging into the so-called “neutrino floor/fog” [12, 13].
@@
RELICS is a reactor neutrino experiment using the LXeTPC technology. LXeTPC is one of the leading technologies for rare event searches, such as detecting dark matter and neutrinoless double beta decay [11, 25, 26, 27, 28, 29]. Benefiting from its ultra-low threshold by using ionization-only channel [30, 31] and reduced background, the LXeTPC is an excellent approach for reactor CENS detection. The RELICS experiment is proposed to be located near the Sanmen Nuclear Power Plant in Zhejiang province, China. At a 25 m baseline to the 3 GW reactor core, the average neutrino flux available for RELICS is 1013cm-2s-1.
@@ IV Expected CENS events
where is Fermi’s coupling constant, is the weak nuclear charge, is the ground state elastic form factor, is the mass of the target nucleus, is the nuclear recoil energy while the incident neutrino energy. is required to ensure coherence, where is the nuclear radius. Normally, this can be satisfied when MeV.
@@ V Background contributions
single-scatter: Given the small interaction cross-section of neutrino, the probability of a neutrino having more than one interaction in the sensitive LXe volume is negligible. We required that the Eer must be less than 0.05 keVer in an NR event, and the second largest energy deposition of Eer less than 5% of the largest one in an ER event.
@@
In addition, dedicated selection methods are developed to suppress backgrounds out of the delayed electrons, such as waveform-classifier and pattern-classifier. These two criteria are described in detail in (V.5).
@@ V.1 Background induced by cosmic-ray neutrons
The cosmic rays that travel through the atmosphere can produce a variety of radiation particles, such as neutrons, protons, pions, etc., through spallation reaction on nitrogen, oxygen, etc [48, 49]. Among them, the high-energy cosmic-ray neutrons (CRNs) can easily pass through the shield and deposit energy in the RELICS detector. The main purpose of the water shield design is to block the CRNs. The size of the water tank is optimized based on the MC simulation. The CRN flux is calculated using the Cosmic-Ray Shower Generator (CRY) source [50]. The energy spectrum of the CRNs from CRY is shown in figure 6. Analyze the MC data shows that only CRNs with energy greater than 103 MeV can deposit energy in LXe, which indicates that the neutrons from the nuclear reactor and laboratory environment normally with energy less than 10 MeV [51] that also shows in figure 6, can not contribute to the backgrounds.
After processing the MC data, and passing the LXe-veto, single-scatter and FV selections, the CRNs-induced NR background rate is estimated to be (7.7 0.7) in keVnr region-of-interest (ROI) of CENS searching, with the recoil energy spectrum shows in figure 7 (top). In addition, the secondaries of CRNs will also induce ER background in LXe, with a rate of (3.9 0.3), which is shown in figure 7 (bottom).
@@ V.2 Background induced by cosmic muons
The RELICS detector will operate without overburden. Hence, special attention has to be paid to the muon-related background. Even though muons can be tagged with 99% efficiency by the PS veto detector, the neutron background produced by muon interacting with the detector materials is still one of the main background components. Dedicated simulations are performed to evaluate the muon-induced backgrounds. The energy and angular distributions of muons follow the Shukla model [52]. The energy deposition in LXe that is caused by muon-induced fast neutron and related gammas are recorded by RelicsSim, and then filtered with LXe-veto, single-scatter and FV selections. Assuming a 99% muon tagging efficiency in 300 veto time, the muon-induced NRs and ERs are (0.08 0.01) in keVnr and (28.0 2) in keVer, respectively. The corresponding spectra are shown in figure 7.
@@ V.3 Neutrons and gammas from the detector materials
Materials constructing the detector contain radioactive contamination. The selection of the materials requires an extensive radioactivity screening campaign for rare event search experiments. The radioactive contamination for all the materials is considered in the RELICS MC model to evaluate the background contribution from these materials. Currently, the radioactivity of isotopes from XENON100 [53] is adopted. A dedicated effort is also being carried out to select clean materials for the RELICS experiment, leading to similar radioactivity for most of the materials.
The NR background rate induced by neutrons is around one event per year and completely negligible in the CENS ROI. ER background induced by gammas is studied in this work. The decay of these isotopes is generated and confined uniformly in the corresponding detector component in the simulation. By applying successive selection criteria, including LXe-veto, single-scatter and FV selections, the spatial distribution of ER background events inside the active LXe volume in the energy region of [0, 20] keVer is shown in figure 8. Figure 7 (bottom) shows the corresponding energy spectrum, which is more than one order of magnitude higher than other ER background components. The average ER background rate is (310 10) in the FV volume.
S2 width selection: As shown in figure 8, a large amount of background events are located close to the liquid-gas interface. These backgrounds are mainly from the radioactivities of PMTs and can not be suppressed without the selection in “z” in an “S2-only” analysis. As demonstrated in [30], the pulse shape of S2 signals can discriminate background happening at the liquid-gas interface to signals distributed uniformly in depth “z”. A toy MC was conducted to consider the drift and diffusion properties of the electron clouds in LXe to optimize such a selection. We simulate S2s and their pulse shape parameters with “z” distributions shown in figure 8 with energy in the CENS ROI. The S2s are simulated via RelicsAPT where the NEST model is incorporated. The S2 pulse shape is simulated using the electron diffusion model , where is the longitudinal diffusion coefficient, is the electron drift velocity, is the width of single electron waveform [54]. In our simulation. = 16 cm2/s, = 0.174 cm/s, and = 0.19s.
@@ V.4.1 Intrinsic 222Rn and 85Kr
Krypton mixes into xenon gas when it is extracted from the air. 85Kr, regardless of its low abundance of in natural krypton, is one of the main backgrounds in any LXe experiments searching for rare events. The concentration of krypton in xenon is usually reduced below ppt level in dark matter search experiments, such as XENON1T and PandaX-4T, through distillations with dedicated columns [55, 56]. In RELICS, we assume a natKr/Xe concentration of 10 ppt and expect (0.28 0.01)10 as shown in figure 7 (bottom), which is a subdominant component.
The other intrinsic background in LXe comes from the decays of 222Rn, which emanates from any material touching xenon gas or liquid. The emanated 222Rn diffuses homogeneously inside the LXe volume within its relatively long half-life of 3.8 days. The decay of 214Pb is another important background component in the CENS ROI in the decay chain of 222Rn. With similar size of TPC and materials, the radioactivity of 222Rn in RELICS is comparable to XENON100. Assuming a conservative 222Rn radioactivity of 40 as achieved by the XENON100 experiment [53] (20 ), the expected ER background rate is estimated as of (0.96 0.03)10-3 kg-1day-1keV-1 as shown in figure 7 (bottom), which is also a subdominant component in the CENS search.
@@ V.4.2 Cosmogenic isotopes and
Natural xenon does not contain . can be produced from when cosmogenic muons kick off a neutron in the nucleus with a cross-section of (2.74 0.4) cm2. Additionally, we consider another production channel through the neutron capture on with an upper limit cross-section of 0.7cm2 at 95% C.L [58]. The production rate of in LXe is calculated based on its production cross-section, the muon’s fluxes and energy spectrum, and thermal neutrons. The decays to via electron capture (EC) process, with a half-life of 36.4 days. We assume that the production and decay of reach equilibrium before deploying the RELICS detector inside the shielding. The estimated equilibrium decay rate is 9.29 atoms/day/kg. The shielding system of RELICS can attenuate neutrons efficiently, and the equilibrium decay rate can decrease to 0.79 atoms/day/kg after 300 days.
@@
We consider three possible mechanisms for production in xenon, following the approaches in [59]. Firstly, the extremely low content of in the atmosphere, which is generated by spallation of through (n, 4n), neutron capture of , or cosmic bombardment of calcium-containing soils, via (n, ). This mechanism only needs to be considered in the context of air leaks of an LXe detector, which is strictly limited to below 0.1 liter per year to ensure the purity level of LXe. Secondly, the spallation of detector materials, especially iron in steel. However, its contribution is greatly suppressed due to the slow diffusion rate of argon in steel. Lastly, the direct spallation of xenon atoms by high-energy cosmogenic protons and neutrons. Only the last possibility is considered in this work. The semi-empirical formula by Silberberg and Tsao was used to calculate the spallation cross-section. The CRY is used to calculate the proton spectrum. Gordon’s neutron spectrum [60] is adopted for cosmogenic neutron estimation. The production rate of this process is estimated to be 0.024 atoms/kg/day with an uncertainty of 100%. The produced decays to the ground state of via EC and subsequently deposits energy via x-rays or Auger electrons. The total energy (relative intensity) for K, L, and M shells are 2.82 keV (90.2%), 0.270 keV (8.9%), and 0.018 keV (0.9%), respectively. As a results, the background rate of in the CENS ROI is below kg-1day-1 and completely negligible for CENS search.
@@ V.5 Delayed electrons
The first kind of delayed electrons is produced via photo-ionization of impurities or metals in the drift region of the TPC. Photo-ionized delay electrons only appear shortly (within a time range that is comparable to the largest drift time of the LXeTPC) after the high energy S2 signals and will be completely subdominant when a 1 exposure is rejected for each of the high energy S2s, leading to an exposure loss of 1%.
Besides photo-ionization, delayed electrons can also be produced in a larger time range of up to a few seconds. The rate of delayed electrons is observed to be correlated with the level of impurities in LXe and the amplitude of the extraction field. Muons passing through the LXeTPC result in large energy deposition, generating a substantial number of ionization electrons. These delayed electrons may occur after the primary S2 signal of muon events, leading to a prolonged tail of signals composed of a single or a few electrons. Due to the high rate of cosmic muons, RELICS faces a major challenge in suppressing background from accidental pile-ups of delayed electrons in a short time range that is comparable to the width of a physical S2 signal. This background is hereafter referred to as Delayed Electron (DE).
@@ V.5.1 Expected rate of DE
As shown in previous studies with XENON1T [61] and LUX [62], the emission rate of delayed electrons are correlated, in time and position, with the previous high-energy events. This part of the background is referred to as a correlated delayed electron background and can be significantly suppressed due to its unique correlation with proceeding high-energy events. We simulate muons passing through the TPC active volume using Geant4, tracking their interaction vertex and the energy deposited along the track. The ionization signals are generated along the muon track, assuming a constant ionization yield of 50 electrons per keV. We then simulated the generation of delayed electrons, assuming that 0.1% of the total ionization electrons from primary interactions are delayed emissions in a time range of 0.1 ms to 2 s, comparable to [61]. The distribution in “delayed time” (time respected to the primary S2) for delayed electrons is assumed to follow a power law with a coefficient of -1.1, as shown in XENON1T [61]. These delayed electrons’ horizontal positions () are assumed to be the same as their primary S2s. Position reconstruction resolution for single electron S2s, which is expected to be 20 mm from optical simulations, is taken into account in this analysis. Ionization S2 signals are grouped together if the time gap between them is below 1 s in order not to split S2 waveforms of events happening at the bottom part of the TPC. As a result, the pile-up of DE (called Pile-up DEs) leads to a significant background in CENS search. In RELICS, we expect more than 10 Hz of muon events in the LXeTPC, leaving a large amount of single electron S2 signals. The expected event rate of a single-electron pile-up background is summarized in Tab. 2. Without implementing appropriate cuts to reduce the rate of these pile-up events, CENS signals with 3 to 6 electrons will be completely overwhelmed.